# Referenced internal-line double copy and application to gauge and gravitational beta functions

### v1

@@ Abstract
Following the approach of Refs. Tourkine:2012vx ; Geyer:2015jch , the double-copy-like decomposition of exchanged internal states in the world-line limit of one-loop string amplitudes is systematically formulated and generalized to both bosonic and heterotic string theories. As an application, the one-loop beta functions for the gauge and gravitational coupling constants are investigated by analyzing the low-energy field-theory limit of the corresponding three-point one-loop amplitudes in heterotic string theory under a naive $T^{6}$ compactification. Due to supersymmetry, these beta functions vanish trivially. However, by decomposing the scattering integrand according to the different internal loop-exchanged states, the most general model-independent results are obtained.
@@ 1 Introduction
The double copy relation reveals a profound connection between the perturbative scattering amplitudes of gauge theories and gravitational theories, expressed as $\text{GR}=\text{YM}^{2}$ Bern:2019prr . Although the tree-level double copy originates from the KLT relations Kawai:1985xq in string theory, recent advances in quantum field theory—such as the 5-loop calculation Bern:2018jmv and the classical solution double copy Monteiro:2014cda —have extended beyond the current reach of perturbative string theory. Nevertheless, string amplitudes continue to serve as an important laboratory for studying scattering amplitudes.
The main focus of research on the perturbative double copy has traditionally been on the correspondence between external states. However, studies about low energy limit of one loop type II string amplitude in Ref.Tourkine:2012vx and similar work on ambitwistor string Geyer:2015jch have opened new avenues for the double copy structure of internal states exchanged at one loop. [^1] Such decomposition is also extended to the two loop amplitude of ambitwistor string.Geyer:2019hnn . Extending these ideas to the bosonic and heterotic string theories forms the central theme of this paper. Furthermore, by adopting the chiral splitting effective formalismDHoker:1988pdl ; DHoker:1989cxq , one-loop integrands of closed string have manifest double copy relation between the left and right movers, which is extensively used in this paper.
@@
where $g$ is the gauge coupling constant, $n_{X}$ and $C_{X}$ with $X=v,f,s$ denotes the numbers and the corresponding quadratic Casimir operators of the representations of the relevant vectors (spin-$1$), fermions (spin-$\frac{1}{2}$) and scalars (spin-0) respectively. The non-Abelian nature of QCD gauge theory results in a negative beta function, leading to the celebrated asymptotic freedom of the running gauge coupling constant. This serves as an ideal test case to examine whether the double copy of the reference internal line yields a correct result.
Moreover, given the non-Abelian nature of general relativity (GR), it is intriguing to investigate the loop corrections to the gravitational beta function arising from the graviton (spin-2), dilaton (spin-0, scalar), antisymmetric tensor field (spin-0, axial scalar in 4D), and gravitino (spin-3/2). At the same time, one may expect that gravity also induces corrections to the gauge beta function, proportional to the gravitational constant $\kappa$, where $\kappa^{2}=32\pi G$ and $G$ is the Newtonian gravitational constant. Since the gauge coupling beta function in four dimensions is dimensionless, one would naively expect the following general form:
@@
where $a$ is a coefficient to be determined by one-loop calculations. A nonvanishing coefficient $a$ would imply a power-law running of the gauge coupling constant due to quantum gravitational effects. The possibility of such gravitational power-law corrections in Eq. (2) was first proposed by Robinson and Wilczek Robinson:2005fj , and had resulted in a long time debate Pietrykowski:2006xy ; Toms:2007sk ; Ebert:2007gf ; He:2010mt ; Anber:2010uj ; Ellis:2010rw . This issue is also addressed in the present work with referenced internal line double copy.
Several works in the literature Minahan:1987ha ; DiVecchia:1996uq ; DiVecchia:1996iz derived the gauge beta function from the low-energy limit of string amplitudes. However, these calculations are restricted to specialized string models with fixed field content $\{n_{v},n_{s},n_{f}\}$ . In contrast, the referenced internal-line double-copy approach delivers a significant advantage: it directly yields a universal result valid for arbitrary $\{n_{v},n_{s},n_{f}\}$.
@@ 2.1.1 Structure of one loop amplitude
and $f(z)$ ($\bar{f}(\bar{z})$) is a function of $\partial^{1,2}G(z)$ and momentum $k_{i}$ and polarization vectors $\zeta_{i}$ ($\xi_{i}$) of external states. One can see that the double copy structure is restored after introducing the loop momentum $l$ with the cost of non-manifest modular symmetry. More serious loop level KLT relation Stieberger:2023nol was also formulated under this formalism.
@@ 2.2.2 Scattering amplitude with odd spin structure
which reflects the fact that only massless fermions are exchanged in these diagrams as well as quantum field theory.
@@ 3.1 Three-point One-Loop Amplitude of Heterotic String
It is obvious to see that $\Lambda^{b}(z)$ is the complex conjugate of $\Lambda^{1}(\bar{z})$.
@@ 3.3.1 Non-pinched diagram
Before evaluating the integral, we observe that not all terms in the above equations contribute to the renormalization of the three-gluon tree-level vertex. However, it is nontrivial to determine whether the $l$-dependent terms affect the physical running. Therefore, it is more appropriate to study the typical integrals over the loop momentum and the fundamental region of $\tau$.
@@ 3.4 Gravitational beta function
In conclusion, perturbative gravitational renormalization is absent in string models compactified from the heterotic string,
@@
3.5 Remark on the gravitational correction to gauge beta function
@@
Then Eq.101,Eq.102 and Eq.103 shows that $\partial G(z)$ and $\partial^{2}G(z)$ have no contribution except $G(z)$. Especially, non-trivial part of $I_{L}^{\text{vector}}$ comes from the $q$-expansion of $\mathcal{J}_{3}$. One instantly realize that
@@
    For non-pinched diagram, the momentum conservation and on-shell condition cause
@@
The gravitational sector, including the graviton, antisymmetric tensor, and dilaton, does indeed receive non-zero quantum corrections to the three-gluon one-loop diagram. However, as argued in Section 3.4 regarding the renormalization conditions for the gravitational beta function, this loop correction generates higher-dimensional operators rather than renormalizing the gauge coupling. Such that
@@
Thanks to the internal-line double copy framework, our analysis and calculations verify these results across a broad class of theories derived from heterotic string theory, in contrast to case-by-case Feynman diagram computations. This result also relies on the perturbative approach and the non-perturbative correction from gravity to gauge coupling remains unknown.
@@ 4 Summary
In Sec.3, we applied the referenced internal line double copy to compute the physical beta functions for the gauge coupling constant and the gravitational coupling constant. The detailed scattering integrands were presented in Sec.3.1. Thanks to the chiral-splitting effective formalism, the double copy structure of the integrands became manifest. One can observe that the difference between these two integrands lies in their left-moving parts, which are $\Phi_{\alpha_{L}\beta_{L}}$ and $\Lambda^{b}$, respectively. In the quantum field theory computation of the beta functions, multiple Feynman diagrams contribute. These diagrams are realized in the string framework through different regions of moduli space, and the corresponding decomposition was discussed in Sec.3.2.
@@ Acknowledgements.
I sincerely thanks Hong-Jian He for discussing the gravitational correction to gauge coupling constant.

### v2

# Referenced internal-line decomposition of one-loop string integrand and application to gauge and gravitational beta functions

@@ Abstract
Following the idea of Refs. Tourkine:2012vx ; Geyer:2015jch , the double-copy-like decomposition of one loop string integrand according to the loop-exchanged internal states is systematically reformulated and generalized to both bosonic and heterotic string theories. As an application, the one-loop beta functions for the gauge and gravitational coupling constants are investigated by analyzing the low-energy field-theory limit of the corresponding three-point one-loop amplitudes in heterotic string theory under a naive $T^{6}$ compactification. Due to supersymmetry, these beta functions vanish trivially. However, by decomposing the scattering integrand according to the different internal loop-exchanged states, the model-independent results of these beta functions are obtained.
@@ 1 Introduction
The double copy relation reveals a profound connection between the perturbative scattering amplitudes of gauge theories and gravitational theories, expressed as $\text{GR}=\text{YM}^{2}$. The most typical $L$-loop $n$-point scattering amplitude with manifest double copy structure can be written as
$A^{n}_{L}\sim\int d^{LD}l\sum_{i}\frac{1}{S_{i}}\frac{N_{i}\tilde{N}_{i}}{D_{i}},$ (1)
where $l$ denotes the loop momentum, $i$ ranges among all the possible $\phi^{3}$ Feynman diagrams, $D_{i}$ denotes the products of corresponding scalar propagators, $S_{i}$ denotes possible symmetrical factors, $N_{i}$ and $\tilde{N}_{i}$ are called BCJ numerators which satisfy the color-kinematic dualityBern:2008qj ; Bern:2019prr .
Historically, the tree-level double copy originated from the KLT relations Kawai:1985xq in tree-level string amplitude. While the celebrated CHY formalismCachazo:2013gna ; Cachazo:2013hca ; Cachazo:2013iea extended the tree level double copy relation to other theories. And later ambitwistor stringMason:2013sva ; Geyer:2014fka ; Geyer:2015jch ; Geyer:2019hnn further spelled the loop amplitude as well as the loop level double copy. Besides, it is also possible to explore the loop level scattering integrand by the forward limit of tree level scattering integrandHe:2015yua ; Cachazo:2015aol ; Cao:2025ygu . By the way, one can also study the field theory double copy by low energy limit of string amplitudeHe:2017spx ; Mafra:2017ioj ; Mizera:2019blq ; Casali:2020knc ; Britto:2021prf ; Balli:2024wje . Although recent advances in quantum field theory, such as the 5-loop calculation Bern:2018jmv and the classical solution double copy Monteiro:2014cda , have extended beyond the current reach of perturbative string theory. Nevertheless, string amplitudes continue to serve as an important laboratory for studying scattering amplitudes.
It is noticeable that the information of the loop-exchanged internal states are not manifest in Eq.1. In contraction, the BCJ numerators $N_{i}$ and $\tilde{N}_{i}$ are functions of external states. So we argue that the main focus of research on the perturbative double copy has traditionally been on the correspondence between external onshell states. However, studies about low energy limit of one loop type II string amplitude in Ref.Tourkine:2012vx and similar work on ambitwistor string Geyer:2015jch have opened new avenues for the double copy structure of internal states exchanged at one loop. [^1] Such decomposition is also extended to the two loop amplitude of ambitwistor string.Geyer:2019hnn . Extending these ideas to the bosonic and heterotic string theories forms the central theme of this paper. Furthermore, by adopting the chiral splitting effective formalismDHoker:1988pdl ; DHoker:1989cxq , one-loop integrands of closed string have manifest double copy relation between the left and right movers, which is extensively used in this paper.
@@
where $g$ is the gauge coupling constant, $n_{X}$ and $C_{X}$ with $X=v,f,s$ denotes the numbers and the corresponding quadratic Casimir operators of the representations of the relevant vectors (spin-$1$), fermions (spin-$\frac{1}{2}$) and scalars (spin-0) respectively. The non-Abelian nature of QCD gauge theory results in a negative beta function, leading to the celebrated asymptotic freedom of the running gauge coupling constant. This serves as an ideal test case to examine whether the decomposition of the reference internal line yields a correct result. Several works in the literature Minahan:1987ha ; DiVecchia:1996uq ; DiVecchia:1996iz derived the gauge beta function from the low-energy limit of string amplitudes. However, these calculations are restricted to specialized string models with fixed field content $\{n_{v},n_{s},n_{f}\}$ . In contrast, the referenced internal-line decomposition delivers a significant advantage: it directly yields a universal result valid for arbitrary $\{n_{v},n_{s},n_{f}\}$.
Moreover, given the non-Abelian nature of general relativity (GR), it is intriguing to investigate the loop corrections to the gravitational beta function arising from the graviton (spin-2), dilaton (spin-0, scalar), antisymmetric tensor field (spin-0, axial scalar in 4D) and gravitino (spin-3/2). However, due to the complicated Feynman rules, it is very hard to perform such loop level calculation in quantum field theory. And the loop-exchanged states information is not clear in the traditional loop-level double copy as shown as Eq.1. Fortunately, we shall see that these loop states present as a double copy after using chiral-splitting effective formalism and internal-line decomposition.
At the same time, one may expect that gravity also induces corrections to the gauge beta function, proportional to the gravitational constant $\kappa$, where $\kappa^{2}=32\pi G$ and $G$ is the Newtonian gravitational constant. Since the gauge coupling beta function in four dimensions is dimensionless, one would naively expect the following general form:
@@
where $a$ is a coefficient to be determined by one-loop calculations. A nonvanishing coefficient $a$ would imply a power-law running of the gauge coupling constant due to quantum gravitational effects. The possibility of such gravitational power-law corrections in Eq. (3) was first proposed by Robinson and Wilczek Robinson:2005fj , and had resulted in a long time debate Pietrykowski:2006xy ; Toms:2007sk ; Ebert:2007gf ; He:2010mt ; Anber:2010uj ; Ellis:2010rw . This issue is also addressed in the this work with referenced internal line decomposition and double copy.
@@ 2.1.1 Structure of one loop amplitude
and $f(z)$ ($\bar{f}(\bar{z})$) is a function of $\partial^{1,2}G(z)$ and momentum $k_{i}$ and polarization vectors $\zeta_{i}$ ($\xi_{i}$) of external states. One can see that the double copy structure is restored after introducing the loop momentum $l$ with the cost of non-manifest modular symmetry. The loop level KLT relation Stieberger:2023nol was also formulated under the chiral splitting effective formalism.
@@ 2.2.2 Scattering amplitude with odd spin structure
which reflects the fact that only massless fermions are exchanged in these diagrams as well as quantum field theory. The internal-line decomposition is trivial in odd spin structure.
@@ 3.1 Three-point One-Loop Amplitude of Heterotic String
It is obvious to see that $\Lambda^{b}(z)$ is the complex conjugate of $\Lambda^{1}(\bar{z})$ with replacing $\xi_{i}$ by $\zeta_{i}$.
@@ 3.3.1 Non-pinched diagram
Before evaluating this integral, we observe that not all terms in the above equations contribute to the renormalization of the three-gluon tree-level vertex. However, it is nontrivial to determine whether the $l$-dependent terms affect the physical running. Therefore, it is more appropriate to study the typical integrals over the loop momentum and the fundamental region of $\tau$.
@@ 3.4 Gravitational beta function
In conclusion, perturbative gravitational renormalization is absent in any string model compactified from the heterotic string,
@@
3.5 The gravitational correction to gauge beta function
@@
Then Eq.102,Eq.103 and Eq.104 shows that $\partial G(z)$ and $\partial^{2}G(z)$ have no contribution except $G(z)$. Hence the non-trivial part of $I_{L}^{\text{vector}}$ only comes from the $q$-expansion of $\mathcal{J}_{3}$. One instantly realize that
@@
    For non-pinched diagram, the momentum conservation and on-shell conditions lead to
@@
It reals that the gravitational sector, including the graviton, antisymmetric tensor, and dilaton, does indeed provide non-zero quantum corrections to the three-gluon one-loop diagram if some supersymmetries are broken. However, as argued in Section 3.4 regarding the renormalization conditions for the gravitational beta function, this loop correction generates higher-dimensional operators rather than renormalizing the gauge coupling. Such that there always is
@@
Thanks to the internal-line decomposition and double copy framework, our analysis and calculations verify these results across a broad class of theories derived from heterotic string theory, in contrast to case-by-case Feynman diagram computations. This result also relies on the perturbative approach and the non-perturbative correction from gravity to gauge coupling remains unknown.
@@ 4 Summary
In Sec.3, we applied the referenced internal line decomposition and double copy to compute the physical beta functions for the gauge coupling constant and the gravitational coupling constant. The detailed scattering integrands were presented in Sec.3.1. Thanks to the chiral-splitting effective formalism, the double copy structure of the integrands became manifest. One can observe that the difference between these two integrands lies in their left-moving parts, which are $\Phi_{\alpha_{L}\beta_{L}}$ and $\Lambda^{b}$, respectively. In the quantum field theory computation of the beta functions, multiple Feynman diagrams contribute. These diagrams are realized in the string framework through different regions of moduli space, and the corresponding decomposition was discussed in Sec.3.2.
@@ Acknowledgements.
I sincerely thank Hong-Jian He for suggesting and discussing the problem about loop correction to gauge and gravitational beta functions. And the two figures in this paper are recycled from the unpublished draft with Hong-Jian He on string loop correction to gauge beta functions by traditional RNS formalism without internal-line decomposition and chiral splitting effective formalism.